# Generator-Based Fuzzers with Type-Based Targeted Mutation

### v2

@@ 3.2.2 Type-Based Mutation
In this step, we use the dynamically captured type distances maintained in $\mathcal{K}$ to select a type to mutate. The probability of choosing a type is inversely proportional to its distance. When a type is selected we then match it up with EI objects within the FCI. If an EI object is found with a matched type from the previous step, SpotOn proceeds in two steps. First, SpotOn decreases the type’s distance (multiplying it by 3/4). Second, SpotOn randomly mutates a subsequence of a random length starting from the matched EI object. If multiple EI objects were matched, then we randomly chose EI objects to mutate their values. If no EI object was found with a matching type, then the type’s distance is updated (multiplying it by 4/3) to decrease the type’s probability of being selected in future iterations.
@@
Table 1: List of AWS lambda applications obtained from GitHub with the number of enclosed lines of code (loc)
@@
4 Evaluation
Our research questions are:
RQ1: Does the SpotOn achieve more coverage than GBF?
@@ 4 Evaluation
RQ2: What is the effectiveness of the type-based targeting without the constant string optimization?
RQ3: What is the overhead of the type-based targeted mutation?
@@
We evaluated our technique using AWS Lambda applications. These applications tend to be compact in the size of user-written code while using extensive code from third-party libraries. Additionally, their inputs often possess a degree of complexity and structure, albeit loosely defined. We collected our benchmark suite from GitHub. We searched for AWS Java Lambdas that are triggered by an S3EventNotification and that use a single S3, and/or a single DynamoDB service. We removed applications with no branches and those that use unsupported formats.
Tab. 1 lists of benchmarks we used, which include:
@@
1) s3-java [25]: when an image file is inserted/updated, the lambda scales it and puts it back in the bucket.
@@
2) anishsana [26]: when a CSV file of students’ grades is inserted/updated into an S3 bucket, the lambda computes an average grade and places a record into a DynamoDB table.
@@
3) load-historic-data [27]: when a stock ticker CSV file, the lambda collects some data from the file and creates a corresponding record onto a DynamoDB table.
@@
4) lambda-unzip [28]: an archive is inserted/updated into an S3 bucket, the lambda unzips its content into the bucket and removes the archived file.
@@
5) csv-loader [29]: when an archived CSV file is inserted/updated into an S3 bucket, the lambda it extracts the CSV file; and creates a DynamoDB entry for each record within the CSV file.
@@
6) nikoshen [30]: when an object is inserted/updated into an S3 bucket, the lambda it checks whether a matching DynmoDB item exists. If not, it creates it.
@@
7) upload-survey [31]: when an object is inserted/updated into an S3 bucket, the lambda serializes the received data, and updates the DynamoDB with their information.
@@
-Implementation: We used Zest [10, 32], the state-of-the-art Java fuzzing tool to add our extension. Zest is a GBF that allows users to write input generators. To compute coverage, Zest instruments dynamically loaded classes to collect coverage information. We built our extension on top of Zest; we call our tool SpotOn. Our implementation utilizes an experimental codebase within Zest that defines FCI as a set of execution indices with values. We extended this work to amend the types of the generator’s call stack to each execution index object of the FCI. We managed child generation by implementing a heuristic in both tools, ensuring inputs with significant time aren’t overrun with excessive children. As processing time for children increases, the number of tests decreases. This heuristic embodies numCandidates$(i_{c})$ in Alg. 1. Additionally, we utilized the Wala [17] framework for static analysis, running it offline before executing SpotOn. Lastly, in SpotOn, mutation alternates between targeted and random mutations with a 50% probability.
@@
Table 2: Average application coverage, (+SpotOn): only found only by SpotOn  (+Zest): only found by Zest. () indicates statically significant results (p-value = 0.05)
@@
-Generators: To fuzz AWSLambda applications, we wrote an AWS generator package, which populates AWS resources interacting with the AWSLambda application, such as S3 and DynamoDB instances. We assume that the structure of the DynamoDB table is known; i.e., its primary key and range key; if one exists. Before every test, the entire AWS state is cleared and repopulated. We randomly generate files of various types and add them to an S3 bucket. Supported types include XML, image, CSV, text, Java class, JavaScript, and archive files. Some of these formats already existed in Zest, others were created by the authors of this paper. These generators have a 0.003 probability of creating erroneous files, like missing or incorrect file extensions. In contrast, for DynamoDB, we randomly generate items with various column sizes (excluding primary and range keys) and populate the resource with them.
-Setup: We used LocalStack 1.4.0 [33]: a local cloud software development framework used to develop, test and run AWS applications locally. We ran our experiment over a Dell Inc. Precision T3600 machine with 32 GB RAM running Ubuntu 22.04.4 LTS. In our evaluation, we utilized Zest’s Maven plugin, JQF, which instruments coverage code within dynamically loaded classes for fuzzing. Coverage of already loaded classes (e.g., java.lang, etc.) is not included in the results of this section. Additionally, coverage for infrastructure applications like com.amazonaws and com.ibm.wala was excluded, as it is beyond the scope of this research.
Table 3: Average of all (third-pary-included) coverage, (+SpotOn): only found by SpotOn  (+Zest): only found by Zest. () indicates statically significant results (p-value = 0.05)
@@
4.1 RQ1: Does SpotOn achieve more coverage?
To answer this question, we fuzzed each benchmark using both Zest and SpotOn, for an hour with 20 repetitions. We categorize coverage results into two categories. (1) Application code coverage: defined as code targets (branches) that exist within the parent package of the program under test. (2) All code coverage: defined as all code targets including those in third-party libraries.
Fig. 2 shows results for the application coverage, where Common lists the common code targets found by both tools, while additional ones found by SpotOn, and Zest are listed in +SpotOn, and +Zest, respectively. Fig. 3 shows results for all-code coverage. We notice that in both tables, most of the code targets were found by both tools, though SpotOn on average finds more code targets. This is mostly apparent in nikoshen and upload-survey. In nikoshen type-targeting was useful to focus the mutation on the DynamoDB Item to create a matching and a valid item. In upload-survey type-targeting was useful to focus on the name of the S3 object. Both benchmarks used the constant string lookup to generate specific string constants. In other benchmarks, we notice that type-targeting is useful for more stable results (more in the following section).
@@ 4.1 RQ1: Does SpotOn achieve more coverage?
4.2 RQ2: What is the effectiveness of type-based targeting?
To isolate the effect of the type-based targeted mutation form the constant string optimization, we separated the later into a separate mode: ZestStrOpt. Fig. 10 through Fig. 11 shows the overall coverage for Zest, ZestStrOpt, and SpotOn. The solid line in the graph describes the median for all 20 repetitions of the experiment. While the shaded lines describe the 95% confidence interval.
The first observation is that in all benchmarks, SpotOn performed either the same or better than Zest. In the case of s3-java, neither ZestStrOpt nor SpotOn demonstrate significant coverage improvements over the baseline Zest. However, we believe this outcome inaccurately represents the effectiveness of type-based targeting, primarily due to a limitation in collecting coverage from dynamically loaded libraries. Particularly in this benchmark, the new coverage achieved by type-based mutation stemmed from pre-loaded classes, which were consequently omitted from the overall coverage collection. In lambda-unzip 10, load-historic-data 10, and anishsana 10, we do not see a significant coverage difference. However, we can observe that SpotOn results are more reliable among the three tools in load-historic-data and anishsana. This is apparent due to its relatively smaller confidence interval. Similarly, we can see stable results for SpotOn in nikoshen, csv-loader and csv-loader.
We see significantly better performance for SpotOn and ZestStrOpt in nikoshen 10, and in csv-loader 10. In fact, the median of the baseline Zest is at the very bottom suggesting that most of the runs were performing poorly compared to SpotOn and ZestStrOpt. The orange shading in csv-loader suggests that a run or two from the population, Zest was able to craft the right input that achieved the desired coverage. In nikoshen the string constant was needed to create a matching state between the DynamoDB item and the S3 object. In csv-loader the string constant was needed to create an S3 object with a particular name. In both benchmarks, the type-based targeted mutation was useful to speed up the coverage by focusing the mutation on the substructure with the matching influencing type. In general, we can conclude that SpotOn has the best reliable performance in coverage finding among the three extensions.
@@ 4.2 RQ2: What is the effectiveness of type-based targeting?
Figure 11: Median coverage of upload-survey
4.3 RQ3: What is the overhead of SpotOn?
To identify the performance bottleneck for SpotOn, we divided the fuzzing operations into 4 processes: (1) mutation: the process of mutating subsequences of the FCI, (2) generation: the process of constructing an input by running the input generators on the FCI, (3) testing: the process of running the program under test with the input generated from the generation step, and (4) handling: the process of handling test result, i.e., new coverage, success or failure.
@@ 4.3 RQ3: What is the overhead of SpotOn?
Fig. 12 shows the average generation time per test. We notice that the average generation cost of SpotOn is not too far off from Zest’s. This overhead is due to computing and maintaining the annotated FCI. More precisely, during the generation SpotOn intercept, every method invocation to keep track of EI counters used in their creation. Then, when a java.util.Random method is about to execute, SpotOn creates the EI instance and looks up its corresponding value within the FCI. If found, the generation uses the present value, otherwise a fresh value is created and the triple of the EI object, its annotated type, the fresh value is added to the FCI.
@@
5 Related Work
@@
However, as CGF lacks information about the input structure, its effectiveness can be negatively affected. Grammar-based fuzzers such as CSmith [37], jsfunfuzz [38], and Grammarinator [39] use a declarative form of the input that defines input’s grammar. Some works used the existence of a test corpus to craft new inputs from existing tests [40, 41, 42, 43]. Most of this work relied on understanding grammers and/or features of existing corpus and then use structural mutation to generate new inputs.
Figure 12: Average generation cost per test; errorbars between 25th and 75th percentiles
Generator-based fuzzers (GBF) utilize programmatically defined generators to create structured inputs. We view grammar-based fuzzers and generator-based fuzzers as orthogonal techniques, with the latter naturally imposing a language-type structure on the input. This structural information can be leveraged during input generation, the primary focus of this paper. More recently, Padhye el. [10] introduced generator-based fuzzers by extending generator-based testing [9] to use random input generators. In our work, we extend Padhye el.’s work to enable type-based targeted mutation.
@@
The work we present in this paper falls into the first category of heuristics for picking seeds, which tries to craft a likely useful input using type-based targeted mutation. In contrast with FairFuzz, our work finds the places where mutations are needed, rather than the places that should be avoided. Our intuition is that for any given branch in the program, there are usually a few places in the input that are relevant to reaching the intended branch.
6 Conclusion and Future Work
In this paper, we introduced a novel type-based mutation heuristic with constant string lookup for generator-based fuzzers. Results show improvements in application coverage of an average of 18.2% (geometric mean) over the baseline, and improvements in coverage including third-party code of 43.2% when the set of covered libraries remained the same, and even more so when our approach covers additional libraries.
An essential aspect of our technique involves the composition of generators directly impacting coverage results. In the future, we plan to investigate the automation of their creation. Also, since type-base targeting can be limited when for example, the input is an array, we plan to leverage the object dynamic information to target particular objects of the same type.
7 Data Availability
Code and experiment reproducibility instructions will be made public once the paper is accepted. An anonymized preliminary version is available for review [48].

### v3

@@ III-C2 Type-Based Mutation
In this step, we use the dynamically captured type distances maintained in $\mathcal{K}$ to select a type to mutate. The type selection process works as follows. First, we sort all types based on their distances, then we normalize their distances. Next, we randomly choose a number and return the first type whose distance is greater than or equal to the selected random number. Then, we match the selected type to those in the FCI. If an EI object is found with the type from the previous step, our algorithm proceeds in two steps. First, it decreases the type’s distance (multiplying it by 3/4). Second, it randomly mutates a subsequence of a random length starting from the matched EI object. If no EI object was found with a matching type, then the type’s distance is updated (multiplying it by 4/3) to decrease the type’s probability of being selected in future iterations.
IV Evaluation
Our research questions are:
@@ IV Evaluation
TABLE I: List of AWS lambda applications obtained from GitHub with the number of enclosed lines of code (loc)
@@
RQ1:
Does the SpotOn achieve more coverage than GBF?
RQ2:
@@
What is the effectiveness of the type-based targeting without the constant string optimization?
RQ3:
What is the overhead of the type-based targeted mutation?
@@
We evaluated our technique using AWS Lambda applications. These applications tend to be compact in the size of user-written code while using extensive code from third-party libraries. Additionally, their inputs often possess a degree of complexity and structure, although loosely defined. We collected our benchmark suite from GitHub. We searched for AWS Java Lambdas that are triggered by an S3EventNotification and that use a single S3, and/or a single DynamoDB service. We removed applications with no branches and those that use unsupported formats.
Tab. I lists of benchmarks we used [22], which include:
  1. 1.
@@
    s3-java [23]: when an image file is inserted/updated, the lambda scales it and puts it back in the bucket.
  2. 2.
@@
    anishsana [24]: when a CSV file of students’ grades is inserted/updated into an S3 bucket, the lambda computes an average grade and places a record into a DynamoDB table.
  3. 3.
@@
    load-historic-data [25]: when a stock ticker CSV file, the lambda collects some data from the file and creates a corresponding record onto a DynamoDB table.
  4. 4.
@@
    lambda-unzip [26]: an archive is inserted/updated into an S3 bucket, the lambda unzips its content into the bucket and removes the archived file.
  5. 5.
@@
    csv-loader [27]: when an archived CSV file is inserted/updated into an S3 bucket, the lambda it extracts the CSV file; and creates a DynamoDB entry for each record within the CSV file.
  6. 6.
@@
    nikoshen [28]: when an object is inserted/updated into an S3 bucket, the lambda it checks whether a matching DynmoDB item exists. If not, it creates it.
  7. 7.
@@
    upload-survey [29]: when an object is inserted/updated into an S3 bucket, the lambda serializes the received data, and updates the DynamoDB with their information.
@@
-Implementation: We used Zest [10, 30], the state-of-the-art Java fuzzing tool to add our extension. Zest is a GBF that allows users to write input generators. To compute coverage, Zest instruments dynamically loaded classes to collect coverage information. We built our extension on top of Zest; we call our tool SpotOn [22]. Our implementation utilizes an experimental codebase within Zest that defines the FCI as a set of execution indices with values. We extended this work to amend the types of the generator’s call stack to each execution index object of the FCI. In both tools, we added a heuristic that controls the number of mutants created from each FCI. The heuristic ensures inputs with significant time are not overrun with excessive children. As processing time for children increases, the number of tests decreases. This heuristic embodies numCandidates$(i_{c})$ in Alg. 1. Additionally, we utilized the Wala [15] framework for static analysis, running it offline before executing SpotOn. Lastly, in SpotOn, mutation alternates between targeted and random mutations with a 50% probability.
@@
TABLE II: Average application coverage, (+SpotOn): only found only by SpotOn  (+Zest): only found by Zest. () indicates statically significant results (p-value = 0.05)
@@
-Generators: To fuzz test AWSLambda applications, we wrote an AWS generator package that populates AWS resources interacting with the AWSLambda application, such as S3 and DynamoDB instances. We assume that the structure of the DynamoDB table is known i.e., its primary key and range key; if one exists. Before every test, the entire AWS state is cleared and repopulated. We randomly generate files of various types and add them to an S3 bucket. Supported types include XML, image, CSV, text, Java class, JavaScript, and archive files. Some of these formats already existed in Zest, others were created by the authors of this paper. These generators have a 0.003 probability of creating erroneous files, like missing or incorrect file extensions. In contrast, for DynamoDB, we randomly generate items with various column sizes (excluding primary and range keys) and populate the resource with them.
-Setup: We used LocalStack 1.4.0 [31]: a local cloud software development framework used to develop, test and run AWS applications locally. We ran our experiment over a Dell Inc. Precision T3600 machine with 32 GB RAM running Ubuntu 22.04.4 LTS. In our evaluation, we utilized Zest’s Maven plugin, JQF, which instruments coverage code within dynamically loaded classes for fuzzing. We collected coverage information for condition coverage and method invocation coverage. Coverage of already loaded classes (e.g., java.lang, etc.) is not included in the results of this section. Additionally, coverage for infrastructure applications like com.amazonaws and com.ibm.wala was excluded, as it is beyond the scope of this research.
TABLE III: Average of all (third-pary-included) coverage, (+SpotOn): only found by SpotOn  (+Zest): only found by Zest. () indicates statically significant results (p-value = 0.05)
@@
IV-A RQ1: Does SpotOn achieve more coverage?
To answer this question, we fuzz tested each benchmark using both Zest and SpotOnfor an hour with 20 repetitions. We categorize coverage results into two categories. (1) Application code coverage: defined as coverages of conditional branches and method invocations within the application code. (2) All code coverage: defined as coverages of conditional branches and method invocations within the application code as well as for those happening within all non-excluded libraries.
Tbl. II shows results for the application coverage, where Common lists the common code targets found by both tools, while additional ones found by SpotOn, and Zest are listed in +SpotOn, and +Zest, respectively. The (*) indicates the statistical significance of the coverage, which shows that the results of 4 out of the 7 benchmarks’ results are statistically significant refuting the null hypothesis that the difference in the coverage results between the two tools is only due to random variation.
Tbl. III shows results for all-code coverage, showing that the coverage of three out of the seven benchmarks is statistically significant. In more detail, we notice that in both tables, most of the code targets were found by both tools, though SpotOn on average finds more code targets. This is mostly apparent in nikoshen and upload-survey. In nikoshen type-targeting was useful to focus the mutation on the DynamoDB Item to create a matching and a valid item. In upload-survey type-targeting was useful to focus on the name of the S3 object. Both benchmarks used the constant string lookup to generate specific string constants. In other benchmarks, we notice that type-targeting is useful for more stable results (more in the following section).
@@ IV-A RQ1: Does SpotOn achieve more coverage?
IV-B RQ2: What is the effectiveness of type-based targeting?
To isolate the effect of the type-based targeted mutation from the constant string optimization, we separated the later into a separate mode: ZestStrOpt. Tbl. 15 through Tbl. 15 shows the overall coverage for Zest, ZestStrOpt, and SpotOn. The solid line in the graph describes the median for all 20 repetitions of the experiment. While the shaded lines describe the 95% confidence interval.
The first observation is that in all benchmarks, SpotOn performed either the same or better than Zest. In the case of s3-java, neither ZestStrOpt nor SpotOn demonstrate significant coverage improvements over the baseline Zest. However, we believe this outcome inaccurately represents the effectiveness of type-based targeting, primarily due to a limitation in collecting coverage from dynamically loaded libraries. Particularly in this benchmark, the new coverage achieved by type-based mutation stemmed from pre-loaded classes, which were consequently omitted from the overall coverage collection. This limitation can be circumvented by instrumenting the code prior to the loading of any of these classes, specifically through the use of a Java Agent. Implementing this approach is a task for future work. In lambda-unzip (Tbl. 15), load-historic-data (Tbl. 15), and anishsana (Tbl. 15), while we do not see a significant coverage difference, we can observe that SpotOn results are more reliable among the three tools. This is apparent due to its relatively smaller confidence interval. Similarly, we can see stable results for SpotOn in nikoshen, csv-loader and csv-loader. We see significantly better performance for SpotOn and ZestStrOpt in nikoshen (Tbl. 15), and in csv-loader (Tbl. 15). In fact, the median of the baseline Zest is at the very bottom suggesting that most of the runs were performing poorly compared to SpotOn and ZestStrOpt. The orange shading in csv-loader suggests that a run or two from the population, Zest was able to craft the right input that achieved the desired coverage. In nikoshen the string constant was needed to create a matching state between the DynamoDB item and the S3 object. In csv-loader the string constant was needed to create an S3 object with a particular name. In both benchmarks, the type-based targeted mutation was useful to speed up the coverage by focusing the mutation on the substructure with the matching influencing type. In general, we can conclude that SpotOn has the most reliable performance in coverage finding among the three extensions.
@@ IV-B RQ2: What is the effectiveness of type-based targeting?
Figure 14: Coverage of upload-survey
Figure 15: Avg. gen. per test; errorbars between $25^{th}-75^{th}$ percentiles
IV-C RQ3: What is the overhead of SpotOn?
To identify the performance bottleneck for SpotOn, we divided the fuzzing operations into four processes:
  1. 1.
    mutation: the process of mutating subsequences of FCI,
  2. 2.
    generation: the process of constructing an input by running the input generators on the FCI,
  3. 3.
    testing: the process of running the program under test with the input generated from the generation step,
  4. 4.
    handling: the process of handling test result, i.e., new coverage, success or failure.
@@ IV-C RQ3: What is the overhead of SpotOn?
Tbl. 15 shows the average generation time per test. We notice that the average generation cost of SpotOn is not too far off from Zest’s. This overhead is due to computing and maintaining the annotated FCI. More precisely, during the generation SpotOn intercept, every method invocation to keep track of EI counters used in their creation. Then, when a java.util.Random method is about to execute, SpotOn creates the EI instance and looks up its corresponding value within the FCI. If found, the generation uses the present value, otherwise a fresh value is created and the triple of the EI object, its annotated type, the fresh value is added to the FCI.
@@
V Threat to Validity
The effectiveness of the GBF depends on its generators, and so our results are dependent on the generators that we have constructed. We created these generators because there are no previously written generators for the types we used in our experiments, i.e., AWS-related types. In general, the results can be dependent on what the generators produce and how they are defined. The former is an inherent limitation in GBF, and we addressed it by designing generators to produce random inputs for all decision-relevant data typically used in programs while keeping others, like eventTime and eventVersion in S3Event, constant.
The latter limitation can arise where the linear reuse of values in the FCI may unintentionally alter the structure of created objects (see [32]), potentially impacting Zest’s results. This limitation is mitigated by designing generators that construct objects in a logical order without relying on any specific program under test.
VI Related Work
@@
However, as CGF lacks information about the input structure, its effectiveness can be negatively affected. Grammar-based fuzzers such as CSmith [36], jsfunfuzz [37], Grammarinator [38], and others [39, 40, 41, 42, 43] use a declarative form of the input that defines input’s grammar. Some works used the existence of a test corpus to craft new inputs from existing tests [43, 44, 45, 46]. Most of this work relied on understanding grammers and/or features of existing corpus and then use structural mutation to generate new inputs.
Generator-based fuzzers (GBF) utilize programmatically defined generators to create structured inputs. While both grammar-based fuzzers [39, 47, 42] and generator based fuzzers [10, 48, 49, 50] are more likely to trigger semantic behavior within the program under test, due to their knowledge of the structure of the input, we view them as alternative techniques to create a structure, semantically correct input. However, we observe that the latter naturally imposing language-type structure, which we leverage to focus the generation on likely useful substructures within the input. Padhye el. [10] introduced generator-based fuzzers by extending generator-based testing [9] to use random input generators. In our work, we extend Padhye el.’s work to enable type-based targeted mutation.
@@
The work we present in this paper falls into the first category of heuristics for picking seeds, which tries to craft a likely useful input using type-based targeted mutation. In contrast with FairFuzz, our work finds the places where mutations are needed, rather than the places that should be avoided. More recent work in GBF [49, 32] highlighted the importance of identifying the structure of the input to increase the effectiveness of the mutation. However these techniques focused on orthogonal aspects of mutation, namely, whether change affects the object structure [49], or side-effects of mutations on subsequent choices [32].
VII Future work
One limitation of this approach is its ineffectiveness with recursive types, as the entire FCI will be made up of a few repeated types making the type-based targetting less effective We plan to extend this work to support effective targeting for objects of recursive types.
We focused on testing serverless applications, as they are ideal candidates for our technique. We plan to evaluate our technique on traditional programs. Unfortunately, existing traditional programs rely on using primitive or recursive types, unsuitable for our technique. Thus, we will seek suitable benchmarks for traditional programs to assess our approach.
Finally, we plan to investigate the automatic generation of the generators, a common hurdle for the usability of GBF.
VIII Conclusion
We introduced a novel type-based mutation heuristic with constant string lookup for generator-based fuzzers. Results show improvements in application coverage of an average of 18% over the baseline, and improvements in coverage including third-party code of 43% when the set of covered libraries remained the same, and even more so when our approach covers additional libraries.